Determine the truth value of each of the following statements:/5mrks

(i) Paris is in France or 4+4=8

(ii) 4+4=9 and5+8=11

(iii) m+2is the unit of area or Kg is the unit of mass


(iv) If Kigali is capital city of Rwanda and Kampala is not capital city of Uganda, then


Bujumbura is the capital city of Burundi.


1
Expert's answer
2021-10-21T15:07:25-0400

Solution.

(i) true (logical operation or; "Paris is in France" statement is true and "4+4=8" statement is true )

(ii) false (logical operation and; "4+4=9" statement is false; "5+8=11" statement is false)

(iii) true (logical operation or; "Kg is the unit of mass" statement is true)

(iv) "Kigali is capital city of Rwanda" statement is true; "Kampala is not capital city of Uganda" statement is false. logical operation Kigali is capital city of Rwanda and Kampala is not capital city of Uganda is false. "Bujumbura is the capital city of Burundi" statement is true. The result of the logical operation of implication  "If Kigali is capital city of Rwanda and Kampala is not capital city of Uganda, then Bujumbura is the capital city of Burundi." is true.


Answer. (i) true; (ii) false; (iii) true; (iv) true.